OpenStudy (anonymous):

If x <2, simplify |x - 2| - 4|-6|

OpenStudy (blacksteel):

Since x < 3, we know that (x-2) is going to be negative. Thus we can write |x-2| as -(x-2), or 2 - x. Furthermore, |-6| = 6, so 4|-6| = 24. Thus |x-2| - 4|-6| = 2 - x - 24 = -x - 22

OpenStudy (blacksteel):

*Since x < 2